OpenStudy (anonymous):

Please help!!!1 will give medal!! 7(x+4)=-2(5x-7)+9x

OpenStudy (alexandervonhumboldt2):

7*(x+4)=-2*(5x-7)+9x 7x+28=-10x+14+9x 17x-9x=14-28 8x=-14 x=-14/8=-7/4